# A Physics-Informed Spatiotemporal Deep Learning Framework for Turbulent Systems

### v2

@@ 1 Introduction & Background
Modeling fluid thermodynamics is critical for advancing our understanding of complex physical systems, including atmospheric dynamics (Trenberth & Stepaniak 2003), climate science (Held & Suarez 1994), and industrial processes (Singh & Heldman 2014). Accurate thermodynamic models enable the prediction of phase transitions, heat transfer, and fluid motion, key to optimizing engineering systems and mitigating environmental impacts (Gyftopoulos & Beretta 2012). However, Direct Numerical Simulation (DNS) is computationally expensive due to the wide range of spatial and temporal scales and strong nonlinearities involved (Xu 2015). As a result, statistical surrogate models have emerged as promising alternatives, capable of learning from limited data and enabling longer simulations at significantly lower cost.
Surrogates are effective for modeling turbulence only if they can capture the nonlinear, multiscale structure of the governing dynamics. To this end, reduced-order models such as Proper Orthogonal Decomposition and Koopman operator theory have provided principled, interpretable, and computationally efficient frameworks for simulating dominant flow structures (Rowley 2005, Brunton et al. 2021). More recently, operator learning approaches—such as the Fourier Neural Operator—have shown promise in approximating Partial Differential Equations (PDE) solution operators, scaling to high-resolution turbulent flows (Li et al. 2020). These methods aim to learn solution operators rather than simple input-output maps. However, most are deterministic and lack mechanisms for uncertainty quantification, limiting their reliability in sensitive or chaotic regimes.
An alternative paradigm uses supervised learning to emulate turbulent flows directly from data. Convolutional Neural Networks (CNNs) have been used to predict flow fields from boundary conditions or coarse inputs, while Recurrent Neural Networks (RNNs) model temporal evolution (Guo et al. 2016, Maulik et al. 2021). Convolutional Recurrent Neural Networks (CRNNs) — which combine CNN-based spatial encoding with RNN-based temporal modeling — have shown success in spatiotemporal forecasting of turbulent flows. These models typically employ an encoder to compress spatial structure, an RNN to model dynamics in reduced space, and a decoder to reconstruct high-dimensional output. Applications include two-dimensional cylinder flows (Beiki & Kamali 2023), turbulence reconstruction from sparse data (Fukami & Fukagata 2021), and short-term forecasting in low-turbulence settings (Mohan et al. 2019, Pandey et al. 2022, Straat et al. 2025). These models are however currently limited in that their architectures are suboptimal in long-range forecasting and are not formulated as statistical models, so they don’t allow formal uncertainty quantification.
@@
In this work, we propose a novel Physics-Informed Convolutional Recurrent Neural Network (PI-CRNN) for efficient, long-range forecasting of fluid thermodynamic systems. Our model integrates four key components: (1) spatial feature extraction via a convolutional autoencoder; (2) temporal modeling through a sequence-to-sequence ConvLSTM architecture; (3) physics-informed inference by penalizing violations of mass, momentum, and energy conservation; and (4) uncertainty quantification using conformal prediction (Shi et al. 2015, Bonas et al. 2024). While previous work has explored CRNNs and PINNs separately, to our knowledge, this is the first model that embeds a ConvLSTM within the reduced space of a convolutional autoencoder while enforcing PDE constraints and providing distribution-free prediction intervals. PI-CRNN model is intended to be used as an offline-trained statistical surrogate, which is crucial for a variety of scientific and engineering applications where long time series of a turbulent physical process are needed but DNS is considerably more expensive and time-consuming.
@@
Our testbed is Rayleigh-Bénard convection (RBC), a canonical example in fluid thermodynamics featuring nonlinear heat-driven flow between differentially heated surfaces (Rayleigh 1916). RBC exhibits high-dimensional turbulence and is highly sensitive to initial conditions, making it ideal for evaluating long-term forecast accuracy and uncertainty quantification. While the governing physical equations, i.e., the data-generating process, are deterministic, this sensitivity results in de facto stochastic behavior. We incorporate physical constraints by enforcing the Navier-Stokes equations via penalized optimization (Navier 1822, Stokes 1845), and quantify uncertainty through conformal prediction, which guarantees calibrated prediction intervals under minimal assumptions. The turbulent nature of RBC, coupled with the computational cost of DNS, motivates the development of a statistical surrogate capable of capturing the spatiotemporal dynamics and associated uncertainties at a fraction of the computational expense. PI-CRNN offers a unified, principled framework for surrogate modeling of complex thermodynamic systems, balancing computational efficiency, physical plausibility, and statistical rigor.
@@
The remainder of the manuscript is structured as follows: Section 2 summarizes the turbulent convection data and governing equations; Section 3 introduces the PI-CRNN model; Section 4 presents the physics-informed learning framework; Section 5 discusses uncertainty quantification; Section 6 reports the results; and Section 7 concludes with a discussion of limitations and future directions.
2 Data
@@
This work utilizes simulated data from a two-dimensional RBC flow for dry air, a gravity-driven turbulent flow between two horizontal plates exhibiting a temperature gradient. The dataset, generated via DNS, includes horizontal and vertical velocity components, pressure, and temperature with high spatial and temporal resolution. RBC is governed by the Navier-Stokes equations under the Boussinesq approximation, which encode mass conservation, momentum, and energy conservation for a thermally stratified fluid (Navier 1822, Boussinesq 1903, Rayleigh 1916). The Rayleigh number for this data (a quantity identifying the amount of turbulence in the system) is $\text{Ra = 2.54}\times 10^{8}$. A low Ra (less than $10^{6}$) corresponds to laminar (smooth) flow, while $\text{Ra}=2.54\times 10^{8}$ indicates strongly turbulent convection, making the dataset both scientifically realistic and statistically challenging. This data offers an ideal testbed: the underlying physics are well established, yet the resulting flow is high-dimensional, nonlinear, and turbulent, posing a significant challenge for long-range forecasting. This motivates the development of a surrogate model that can capture both uncertainty and dynamics efficiently. Further technical details on the data and governing equations are in Section S.2 of the supplementary material.
@@
The data was generated on a $n_{x}\times n_{z}$ grid, $n_{x}=n_{z}=256$, with a temporal resolution of $\Delta_{t}=0.1$ seconds for a total of T=2,500T=2{,}500 time steps. We denote the RBC spatiotemporal process with $d=4$ covariates observed on a $n_{x}\times n_{z}$ grid by $\bm{Y}_{t}(x,z)=\left(u_{t}(x,z),w_{t}(x,z),p_{t}(x,z),\theta_{t}(x,z)\right)$, where $u_{t}(x,z)$ and $w_{t}(x,z)$ are the horizontal and vertical velocity components, respectively, $p_{t}(x,z)$ is pressure, $\theta_{t}(x,z)$ is temperature, $(x,z)$ are spatial coordinates, and $t=1,\ldots,T$. For simplicity, we hereto omit the spatial coordinates, so we denote one observation over the entire spatial domain by $\bm{Y}_{t}=\left(u_{t},w_{t},p_{t},\theta_{t}\right)\in\mathbb{R}^{n_{x}\times n_{z}\times d}$. Since the temporal modeling approach we propose considers sequences of data, e.g., $\left(\bm{Y}_{3},\bm{Y}_{4}\right)=f\left(\bm{Y}_{1},\bm{Y}_{2}\right)$”, we denote the sequence of $b$ observations before $t$ by $\mathcal{Y}^{-}_{t;b}=\left(\bm{Y}_{t-b},\ldots,\bm{Y}_{t-2},\bm{Y}_{t-1}\right)$ and the sequence of $a$ observations starting at $t$ by $\mathcal{Y}^{+}_{t;a}=\left(\bm{Y}_{t},\bm{Y}_{t+1},\ldots,\bm{Y}_{t+a-1}\right)$.
@@ 3.1 Spatial Model: Convolutional Autoencoder
The spatial encoder consists of $\ell=1,\ldots,L$ convolutional layers with $K^{\ell}$ kernels in the $\ell$-th layer, where each kernel is a two-dimensional matrix $r^{\ell}_{x}\times r^{\ell}_{z}$ (since the data is observed in two dimensions). The kernels each perform a local convolution of the input every $s=2$ elements (strides) and are summed across the kernels from the $(\ell-1)$-th layer. Strided convolutions combine downsampling and feature extraction in a single operation, reducing computational cost compared to separate convolution and pooling layers (Springenberg & Uhl 2019). As such, the $\ell$-th convolutional layer with $K^{\ell}$ kernels and stride $s$ maps yields $K^{\ell}$ feature maps, each one of dimensions $n_{x}^{\ell}\times n_{z}^{\ell}$, i.e., a three-dimensional matrix $n_{x}^{\ell}\times n_{z}^{\ell}\times K^{\ell}$, where $n_{x}^{\ell}=\lfloor n_{x}^{\ell-1}/s\rfloor$ and $n_{z}^{\ell}=\lfloor n_{z}^{\ell-1}/s\rfloor$. In the $\ell$-th layer, denote by $Z^{\ell}_{i,j,f}$ the $(i,j)$ entry of the $f$-th feature map, by $w^{\ell}_{a,b,\tilde{f},f}$ the $(a,b)$ entry of the $f$-th kernel for the $\tilde{f}-$th channel of the previous layer, $\tilde{f}\in\{1,\ldots,K^{\ell-1}\}$, and by $w^{\ell}_{0,f}$ the bias term. The spatial encoder performs the following:
@@
Figure 1: Architecture of the CAE model. The input $\bm{Y}$ is projected onto reduced space through a series of strided convolutions, (2) in yellow, to $\bm{\tilde{Y}}\in\mathbb{R}^{\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}}$ and subsequently projected back onto the original dimensions, $\mathbb{R}^{n_{x}\times n_{z}\times d}$, through a series of strided deconvolutions, in blue.
@@ 3.2 Temporal Model: Recurrent Neural Network
The temporal structure of $\bm{Y}_{t}$ is modeled in reduced space, i.e., we model the sequence $\tilde{\mathcal{Y}}^{+}_{t;a}=\left(\tilde{\bm{Y}}_{t},\tilde{\bm{Y}}_{t+1},\ldots,\tilde{\bm{Y}}_{t+a-1}\right)$, as a function of the previous $b$ observations $\tilde{\mathcal{Y}}^{-}_{t;b}=\left(\tilde{\bm{Y}}_{t-b},\tilde{\bm{Y}}_{t-b+1},\ldots,\tilde{\bm{Y}}_{t-1}\right)$. The temporal model comprises of two components: a context builder, which learns the temporal structure of the (reduced space) input sequence into a context representation, and a sequence generator, which learns how to generate the future sequence given the context representation. This sequence-to-sequence approach fundamentally differs from autoregressive strategies that predict only a single future timestep at each iteration and then feed that prediction back into the model to forecast subsequent steps. The problem is well-documented in highly nonlinear systems and multistep forecasting, where autoregressive methods tend to diverge from physically realistic trajectories (Pascanu et al. 2013, Sutskever et al. 2014, Taieb & Atiya 2016, Guen & Thome 2019). In Figure S11 of the supplementary material we provide evidence with our data.
@@
Both the context builder and sequence generator are a ConvLSTM model, an RNN that is able to capture short- and long-term temporal dependencies while also accounting for the spatial structure of the data which, in this case, is a sequence of $b$ three-dimensional matrices, each of dimensions $\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}$ (Hochreiter & Schmidhuber 1997, Shi et al. 2015). At time step $t$, a ConvLSTM model takes as input a (reduced space) observation $\tilde{\bm{Y}}_{t}$ along with the hidden and cell state matrices $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, which carry temporal and memory information from previous time steps, and outputs their respective value for the next time step, formally,
@@
Importantly, only the final hidden and cell states, $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, are retained and passed to the sequence generator and the output of each step in the context builder, $\tilde{\bm{Y}}_{t-j+1}$, is discarded. This reflects the context builder’s role to summarize the past into context. A visual representation of the context builder is shown in Figure S2 of the supplementary material.
@@
Importantly, unlike in (7), the output at each step in (8) is fed back into the sequence generator and used as input for the next prediction. This recursive mechanism allows the model to generate multi-step forecasts in a single forward pass while maintaining temporal consistency. A key advantage of this approach is that it avoids the error accumulation and discontinuities inherent in traditional iterative forecasting methods, where entire predicted sequences are repeatedly fed back into the model. Instead, the ConvLSTM’s hidden and cell states propagate sequential dependencies internally, enabling stable long-horizon predictions. A visual representation of the sequence generator is shown in Figure S2 the supplementary material.

### v3

@@ 1 Introduction & Background
Modeling fluid thermodynamics is critical for advancing our understanding of complex physical systems, including atmospheric dynamics (Trenberth and Stepaniak 2003), climate science (Held and Suarez 1994), and industrial processes (Singh and Heldman 2014). Accurate thermodynamic models enable the prediction of phase transitions, heat transfer, and fluid motion—key to optimizing engineering systems and mitigating environmental impacts (Gyftopoulos and Beretta 2012). However, Direct Numerical Simulation (DNS) is computationally expensive due to the wide range of spatial and temporal scales and strong nonlinearities involved (Xu 2015). As a result, statistical surrogate models have emerged as promising alternatives, capable of learning from limited data and enabling longer simulations at significantly lower cost.
Surrogates are effective for modeling turbulence only if they can capture the nonlinear, multiscale structure of the governing dynamics. To this end, reduced-order models such as Proper Orthogonal Decomposition and Koopman operator theory have provided principled, interpretable, and computationally efficient frameworks for simulating dominant flow structures (Rowley 2005, Brunton et al. 2021). Proper orthogonal decomposition-based surrogate modeling combined with Gaussian process emulation has further enabled efficient reduced-order representations with built-in uncertainty quantification by treating reduced states as time-independent (Mak et al. 2018). More recently, operator learning approaches, such as the Fourier Neural Operator, have shown promise in approximating Partial Differential Equations (PDE) solution operators, scaling to high-resolution turbulent flows (Li et al. 2020). These methods aim to learn solution operators rather than simple input-output maps. However, most are deterministic and lack mechanisms for uncertainty quantification, limiting their reliability in sensitive or chaotic regimes.
An alternative paradigm uses supervised learning to emulate turbulent flows directly from data. Convolutional Neural Networks (CNNs) have been used to predict flow fields from boundary conditions or coarse inputs, while Recurrent Neural Networks (RNNs) model temporal evolution (Guo et al. 2016, Maulik et al. 2021). Convolutional Recurrent Neural Networks (CRNNs) — which combine CNN-based spatial encoding with RNN-based temporal modeling — have shown success in spatiotemporal forecasting of turbulent flows. These models typically employ an encoder to compress spatial structure, an RNN to model dynamics in reduced space, and a decoder to reconstruct high-dimensional output. Applications include two-dimensional cylinder flows (Beiki and Kamali 2023), turbulence reconstruction from sparse data (Fukami and Fukagata 2021), and short-term forecasting in low-turbulence settings (Mohan et al. 2019, Pandey et al. 2022, Straat et al. 2025). These models are however currently limited in that their architectures are suboptimal in long-range forecasting and are not formulated as statistical models, so they don’t allow formal uncertainty quantification.
@@
In this work, we propose a novel Physics-Informed Convolutional Recurrent Neural Network (PI-CRNN) for efficient, long-range forecasting of fluid thermodynamic systems. Our model integrates four key components:
  • spatial feature extraction via a convolutional autoencoder
  • temporal modeling through a sequence-to-sequence ConvLSTM architecture
  • physics-informed inference by enforcing mass, momentum, and energy conservation
  • uncertainty quantification using conformal prediction (Shi et al. 2015, Bonas et al. 2024)
@@
While previous work has explored CRNNs and PINNs separately, to our knowledge, this is the first model that embeds a ConvLSTM within the reduced space of a convolutional autoencoder while enforcing PDE constraints and providing distribution-free prediction intervals. PI-CRNN model is intended to be used as an offline-trained statistical surrogate, which is crucial for a variety of scientific and engineering applications where long time series of a turbulent physical process are needed but DNS is considerably more expensive.
@@
Our testbed is Rayleigh-Bénard convection (RBC), a canonical example in fluid thermodynamics featuring nonlinear buoyancy-driven flow between differentially heated surfaces (Rayleigh 1916). RBC exhibits high-dimensional turbulence and is highly sensitive to initial conditions, making it ideal for evaluating long-term forecast accuracy and uncertainty quantification. While the governing physical equations, i.e., the data-generating process, are deterministic, this sensitivity results in de facto stochastic behavior (Lorenz 1963). We incorporate physical constraints by enforcing the Navier-Stokes equations via penalized optimization (Navier 1822, Stokes 1845), and quantify uncertainty through conformal prediction, which guarantees calibrated prediction intervals under minimal assumptions. The turbulent nature of RBC, coupled with the computational cost of DNS, motivates the development of a statistical surrogate capable of capturing the spatiotemporal dynamics and associated uncertainties at a fraction of the computational expense. PI-CRNN offers a unified, principled framework for surrogate modeling of complex thermodynamic systems, balancing computational efficiency, physical plausibility, and statistical rigor.
@@
The remainder of the manuscript is structured as follows: Section 2 summarizes the turbulent convection data and scientific background; Section 3 introduces the PI-CRNN model; Section 4 presents the physics-informed learning framework; Section 5 reports the results; and Section 6 concludes with a discussion of limitations and future directions.
2 Data and Scientific Background
The data employed in this work is numerically simulated two-dimensional Rayleigh-Bénard convection (RBC). RBC is a canonical fluid flow and is a well-studied model of turbulent thermal convection (Lohse and Shishkina 2024). Turbulent thermal convection is a fundamental natural process that governs heat transport in oceanographic, climatic, and astrophysical settings (Grossman and Lohse 2000, Ahlers et al. 2009, Lohse and Xia 2010), and is crucial for HVAC systems and natural draft cooling towers for large-scale nuclear or thermal power stations (Stanford 2003). Due to fluid turbulence (the chaotic and seemingly random evolution of the state of the fluid flow), the physics underlying RBC defies simple analytical treatment. Therefore, understanding and modeling the transport processes in these contexts requires numerical or statistical modeling.
A particularly important application across many natural science and engineering disciplines is that of particle dispersion in turbulent natural convection. Microscopic particles, such as cloud droplets or drizzle, dynamically respond to turbulent thermal convection. Modern approaches to studying particle dispersion often take one of several flavors. For instance, one may study particle dispersion based on artificially prescribed flow statistics (e.g., with Gaussian white noise). The fundamental drawback to this method is that the underlying flow statistics miss the spatial and temporal correlations characteristic of turbulent convection. The result is that the particle response may be fundamentally incorrect. An alternative approach is to numerically solve both the highly non-linear and multi-scale Navier-Stokes equations and the equations of motion for tens of millions of individual particles simultaneously. This method properly represents the statistics of the turbulent convective flow and thus the particle-fluid interactions, but can be computationally expensive and time consuming, thereby rendering ensemble calculations infeasible at large enough scales. Thus, finding an alternative way to efficiently generate ensembles of the turbulent convective flow state, ensuring that underlying statistics are properly represented, is of immense value.
@@ 2 Data and Scientific Background
Importantly, RBC is the mathematical tool by which we study the underlying turbulent processes governing heat and particle transport in the aforementioned contexts. RBC is driven by a gravitational instability due to its density. The density of a fluid is often controlled by its temperature, and for many natural settings, increasing (decreasing) the temperature leads to a decrease (increase) in density. If cool dense fluid is initially found below warmer, less dense, fluid, vertical motion is induced as the system tries to attain a stable state. In the RBC flow considered here, this flow state is driven by a difference in temperature between the warm lower plate and the cold upper plate, which produces convection cells as warm (thus less dense) fluid rises and cold (thus more dense) fluid falls, as seen in Figure S1 of the supplementary material. The dataset, generated via DNS, includes horizontal and vertical velocity components, pressure, and temperature with high spatial and temporal resolution. RBC is governed by the Navier-Stokes equations under the Boussinesq approximation, which encode mass conservation, momentum, and energy conservation for a thermally stratified fluid (Kundu et al. 2024). The Rayleigh number for this data (a quantity identifying the amount of turbulence in the system) is $\text{Ra}=2.54\times 10^{8}$. A low Ra (less than $10^{6}$) corresponds to laminar (non-turbulent) flow, while in our case we have strongly turbulent convection, making the dataset both scientifically realistic and statistically challenging. This data offers an ideal testbed: the underlying physics are well established, yet the resulting flow is high-dimensional, nonlinear, and turbulent, posing a significant challenge for long-range forecasting. This motivates the development of a surrogate model that can capture both uncertainty and dynamics efficiently. Further details on the data and governing equations are in Section S.2 of the supplementary material.
In light of these observations, we address the following research questions:
  • Q1.
    Can a computationally efficient surrogate model reproduce the essential spatial and temporal statistics of strongly turbulent convection?
  • Q2.
    Can such a surrogate enable accurate long-range forecast of turbulent thermal convection?
@@
The data was generated on a $n_{x}\times n_{z}$ grid, $n_{x}=n_{z}=256$, with a temporal resolution of $\Delta_{t}=0.1$ seconds for a total of T=2,500T=2{,}500 time steps. We denote the RBC spatiotemporal process with $d=4$ covariates observed on a $n_{x}\times n_{z}$ grid by $\bm{Y}_{t}(x,z)=\left(u_{t}(x,z),w_{t}(x,z),p_{t}(x,z),\theta_{t}(x,z)\right)$, where $u_{t}(x,z)$ and $w_{t}(x,z)$ are the horizontal and vertical velocity components, respectively, $p_{t}(x,z)$ is pressure, $\theta_{t}(x,z)$ is temperature, $(x,z)$ are spatial coordinates, and $t=1,\ldots,T$. For simplicity, we hereto omit the spatial coordinates, so we denote one observation over the entire spatial domain by $\bm{Y}_{t}=\left(u_{t},w_{t},p_{t},\theta_{t}\right)\in\mathbb{R}^{n_{x}\times n_{z}\times d}$. We also denote the sequence of $b$ observations before $t$ by $\mathcal{Y}^{-}_{t;b}=\left(\bm{Y}_{t-b},\ldots,\bm{Y}_{t-2},\bm{Y}_{t-1}\right)$ and the sequence of $a$ observations starting at $t$ by $\mathcal{Y}^{+}_{t;a}=\left(\bm{Y}_{t},\bm{Y}_{t+1},\ldots,\bm{Y}_{t+a-1}\right)$.
@@ 3.1 Spatial Model: Convolutional Autoencoder
The spatial encoder consists of $\ell=1,\ldots,L$ convolutional layers with $K^{\ell}$ kernels in the $\ell$-th layer, where each kernel is a two-dimensional matrix $r^{\ell}_{x}\times r^{\ell}_{z}$ (since the data is observed in two dimensions). The kernels each perform a local convolution of the input every $s=2$ elements (strides) and are summed across the kernels from the $(\ell-1)$-th layer. Strided convolutions combine downsampling and feature extraction in a single operation, reducing computational cost compared to separate convolution and pooling layers (Springenberg and Uhl 2019). As such, the $\ell$-th convolutional layer with $K^{\ell}$ kernels and stride $s$ maps yields $K^{\ell}$ feature maps, each one of dimensions $n_{x}^{\ell}\times n_{z}^{\ell}$, i.e., a three-dimensional matrix $n_{x}^{\ell}\times n_{z}^{\ell}\times K^{\ell}$, where $n_{x}^{\ell}=\lfloor n_{x}^{\ell-1}/s\rfloor$ and $n_{z}^{\ell}=\lfloor n_{z}^{\ell-1}/s\rfloor$. In the $\ell$-th layer, denote by $Z^{\ell}_{i,j,f}$ the $(i,j)$ entry of the $f$-th feature map, by $w^{\ell}_{a,b,\tilde{f},f}$ the $(a,b)$ entry of the $f$-th kernel for the $\tilde{f}-$th channel of the previous layer, $\tilde{f}\in\{1,\ldots,K^{\ell-1}\}$, and by $w^{\ell}_{0,f}$ the bias term. The spatial encoder performs the following:
@@
Figure 1: Architecture of the CAE model. The input $\bm{Y}$ is projected onto reduced space through a series of convolutions, (2) in yellow, to $\bm{\tilde{Y}}$ and subsequently projected back onto the original dimensions, $n_{x}\times n_{z}\times d$, through a series of deconvolutions, in blue.
@@ 3.2 Temporal Model: Recurrent Neural Network
The temporal structure of $\bm{Y}_{t}$ is modeled in reduced space, i.e., we model the sequence $\tilde{\mathcal{Y}}^{+}_{t;a}=\left(\tilde{\bm{Y}}_{t},\tilde{\bm{Y}}_{t+1},\ldots,\tilde{\bm{Y}}_{t+a-1}\right)$, as a function of the previous $b$ observations $\tilde{\mathcal{Y}}^{-}_{t;b}=\left(\tilde{\bm{Y}}_{t-b},\tilde{\bm{Y}}_{t-b+1},\ldots,\tilde{\bm{Y}}_{t-1}\right)$. The temporal model comprises of two components: a context builder, which learns the temporal structure of the (reduced space) input sequence into a context representation, and a sequence generator, which learns how to generate the future sequence given the context representation. This sequence-to-sequence approach fundamentally differs from autoregressive strategies that predict only a single future timestep at each iteration and then feed that prediction back into the model to forecast subsequent steps. The problem is well-documented in highly nonlinear systems and multistep forecasting, where autoregressive methods tend to diverge from physically realistic trajectories (Pascanu et al. 2013, Sutskever et al. 2014, Taieb and Atiya 2016, Guen and Thome 2019). In Figure S12 of the supplementary material we provide evidence with our data.
@@
Both the context builder and sequence generator are a ConvLSTM model, an RNN that is able to capture short- and long-term temporal dependencies while also accounting for the spatial structure of the data which, in this case, is a sequence of $b$ three-dimensional matrices, each of dimensions $\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}$ (Hochreiter and Schmidhuber 1997, Shi et al. 2015). At time step $t$, a ConvLSTM model takes as input $\tilde{\bm{Y}}_{t}$ along with the hidden and cell state matrices $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, which carry temporal and memory information from previous time steps, and outputs their respective value for the next time step, formally,
@@
Importantly, only the final hidden and cell states, $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, are retained and passed to the sequence generator and the output of each step in the context builder, $\tilde{\bm{Y}}_{t-j+1}$, is discarded. This reflects the context builder’s role to summarize the past into context. A visual representation of (7) is shown in Figure S2 of the supplementary material.
@@
Importantly, unlike in (7), the output at each step in (8) is fed back into the sequence generator and used as input for the next prediction. This recursive mechanism allows the model to generate multi-step forecasts in a single forward pass while maintaining temporal consistency. A key advantage of this approach is that it avoids the error accumulation and discontinuities inherent in traditional iterative forecasting methods, where entire predicted sequences are repeatedly fed back into the model. Instead, the ConvLSTM’s hidden and cell states propagate sequential dependencies internally, enabling stable long-horizon predictions. A visual representation of (8) is shown in Figure S2 the supplementary material.